ELEMENTARY ABELIAN 2-GROUPS ACTING ON BIPLANES (121,16,2) by Vladimir Ćepulić, Zagreb, Croatia

Abstract. The possible actions of elementary abelian 2-groups on hypothetical biplanes (121,16,2) have been investigated and uniquely determined. Mathematical subject classification(1991): 05B05 Keywords: biplane, involution, automorphism group, orbit.